Alexander Corporation reports the following components of stockholders equity at December 31, 2018. Common stock-$25 par valComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Prepare journal entries tRequired 1 Required 2 Required 3 Prepare a statement of retained earnings for the year ended December 31, 2019. (Amounts to bRequired 1 Required 2 Required 3 -------------------- Prepare the stockholders equity section of the companys balance sheet
Answer #1
1
Date General Journal Debit Credit
2-Jan Treasury stock, common 107,500 =4300*25
Cash 107,500
7-Jan Retained earnings 58,050 =(43000-4300)*1.5
Common dividend payable 58,050
28-Feb Common dividend payable 58,050
Cash 58,050
9-Jul Cash 51,600 =1720*30
Treasury stock, common 43,000 =1720*25
Paid-in capital, treasury stock 8,600
27-Aug Cash 43,000 =2150*20
Paid-in capital, treasury stock 8,600
Retained earnings 2,150
Treasury stock, common 53,750 =2150*25
9-Sep Retained earnings 85,140 =(43000-430)*2
Common dividend payable 85,140
22-Oct Common dividend payable 85,140
Cash 85,140
31-Dec Income summary 65,000
Retained earnings 65,000
2
ALEXANDER CORPORATION
Statement of Retained Earnings
For Year Ended December 31, 2019
Retained earnings, December 31, 2018 385,000
Add: Net income 65,000
450,000
Less: Cash dividends declared (143,190)
Less: Treasury stock reissuances (2,150)
Retained earnings, December 31, 2019 304660
3
ALEXANDER CORPORATION
Stockholders’ Equity Section of the Balance Sheet
December 31, 2019
Common stock - $25 par value 1,075,000
Paid-in capital in excess of par value, common stock 86,000
Retained earnings 304,660
Less cost of treasury stock (10,750)
Total stockholders’ equity 1,454,910